### v1

# A call for frugal modelling: two case studies involving molecular spin dynamics

Abstract
As scientists living through a climate emergency, we have a responsibility to lead by example, or to at least be consistent with our understanding of the problem, which in the case of theoreticians involves a frugal approach to modelling. Here we present and critically illustrate this principle. First, we compare two models of very different level of sophistication which nevertheless yield the same qualitative agreement with an experiment involving electric manipulation of molecular spin qubits while presenting a difference in cost of orders of magnitude. As a second stage, an already minimalistic model involving the use of single-ion magnets to implement a network of probabilistic p-bits, programmed in two different programming languages, is shown to present a difference in cost of a factor of . In both examples, the computationally expensive version of the model was the one that was published. As a community, we still have a lot of room for improvement in this direction.
Introduction: doing science during a climate crisis
We live in a society of growth. As a positive result, we have obtained accelerated advances in science and technology that have enabled improved welfare and human development. A negative side result one cannot overlook, however, is anthropogenic climate change, a challenge that is undoubtly deserving of our focused attention. From the technological point of view, innovation and efficiency are being studied as tools to decrease CO emission intensity, and indeed one can expect that this will play an important role in our response to the climate crisis.1, 2, 3, 4 There are also voices that instead call either for limits to growth or even for degrowth, and this point of view will also likely be relevant going forward.5, 6 Here we propose to reconcile both views and put them in the context of modelling of chemical systems of interest in Physics and Materials Science.
The role of science in this crisis has been twofold: quantifying and explaining the processes, and also pointing towards our possible ways out. Abundant datatasets have been employed to estimate key indicators related to forcing of the climate system, including emissions of greenhouse gases and short-lived climate forcers, greenhouse gas concentrations, radiative forcing, surface temperature changes, the Earth’s energy imbalance, warming attributed to human activities, the remaining carbon budget, and estimates of global temperature extremes. Considering this ample variety of aspects, the indicators show that human-induced warming reached 1.14 [0.9 to 1.4] C averaged over the 2013–2022 decade and 1.26 [1.0 to 1.6] C in 2022. Over the 2013–2022 period, human-induced warming has been increasing at an unprecedented rate of over 0.2 C per decade.7 The most relevant decisions towards mitigation and adaptation are political and mostly outside of our hands for those of us working in the academia, and yet an obvious question is: do we need to do any changes in our academic policies and in our day to day work? The first aspect in which the academia has answered ”yes” concerns air travel, and there are some ongoing efforts to reduce our collective carbon footprint in that direction.8 More recently we have also seen increased concerns regarding the carbon footprint from computations, and only in recent times have tools and guidelines been widely available to computational scientists to allow them to estimate their carbon footprint and be more environmentally sustainable.9, 10
Zooming momentarily out from scientific computing to computing in general, computing’s global share of carbon emissions has been estimated to be from as low as 1.8% if one focuses on operating costs to as high as 3.9% if the full supply chain is taken into account.11 More importantly, while most economic sectors overall are starting to design or implement plans to reduce carbon emissions, computing’s emissions are still strongly on the rise. This is despite continuous improvements in efficiency, as these have been consistently overtaken by increases in demand.12 Indeed, emissions from computing, accounting for the production of the devices, have been projected to be close to 80% of our emissions budget by 2040 to limit warming to 1.5ºC.13 This impossibility of continuing with ”business as usual” has resulted in a call for “frugal computing”, where computing is urgently recognized as a limited resource in the sense that it pushes against the hard boundaries of a hospitable planet. Ultimately -but sadly this means sooner rather than later- we need to advance towards zero-carbon computing, whether this means “doing more with less”, or even “doing less with much less”.
These necessary but extremely ample goals need to be translated into each particular context, which in our case means translating frugal computing into frugal modelling. We will need to consider the shared costs, in terms of climate consequences, of our choices when we are doing science. Unfortunately, as a community we have become accustomed to employ increasingly wasteful amounts of resources in our calculations. Fortunately, it is possible to answer plenty of interesting questions in Nanoscience, Chemistry, Physics and Materials modelling while employing a frugal approach. In particular, we will focus here on our own field of expertise, namely magnetic molecules, although the general ideas may be extrapolated to many fields. emissions
Magnetic molecules have been studied for decades, firstly as controllable models for interactions and phenomena in solid-state Physics and more recently molecular nanomagnets have been presented as candidates for bits,14, 15, 16 qubits,17, 18, p-bits,19 and also as components for nanotechnological devices.20, 21, 22, 23 Manipulation of individuals spins, once a distant dream, is today a practical reality, if not one of immediate practical applicability. In parallel, advances have been made in modelling the influence of the chemical environment on said spin states and their dynamics,24, 25, 26, 27, 28, 29, 30 with wildly different computational costs, as we will see below in some detail.
Herein we present a couple of case studies focusing firstly on the contents of the models and then on their implementation. In the next section ”Choosing affordable models” we will present different alternatives for the modelling of the electric field modulation of the ligand field for the coherent control of the spin states in a molecular spin qubits.31 In the section ”Finding inexpensive implementations” we will present a frugal model for magnetic molecules as probabilistic bits,19 which can also serve, less frugally, to model their macroscopic magnetic properties.32 As we will see in these examples, models that can be similarly useful in practice can have costs varying in many orders of magnitude, whereas implementing an already frugal model to a more efficient implementation can significantly further the savings. Note however that unless frugality is maintained as a boundary condition, mere computational efficiency will lead to a rebound effect, increased use (precisely because of the improved return on investment) and increased emissions.
Choosing affordable models
The first questions that define a scientific work are typically ”what is the problem or question we aim to answer?” and, closely related, ”how are we modelling this?”. An informal cost-benefit analysis follows, i.e. with how much resources one needs to invest vs what one gets from it. In current academia, the resources or cost are related with the funding and computational capabilities of the research group, rather than an emissions budget. The benefits, when passed through the academic filter, are still about knowledge and the common good, but projected onto high-profile publications, and currently the refereeing process rewards computational effort, even if it leads to a waste we collectively cannot afford.
To illustrate this problem, let us focus on the different pathways that one can choose to model the effect of the ligand field on the magnetic and spectroscopic properties of metal ions, a question that has received some attention in the past decade in the context of the so-called Single Ion Magnets and molecular spin qubits, since the spin dynamics of magnetic molecules are in large part based in the variation of the energies of the different spin states with distortions of the molecular structure.
The widely accepted standard in this field are ab initio calculations, where CASPT2 or NEVPT2 are considered superior to CASSCF for fundamental reasons, and MOLCAS or ORCA are employed as standard computational codes. A comparatively fringe modelling approach is based on effective charges acting on the f orbitals; this is widely considered much less exact, again for fundamental reasons. It is not often that the predictive power of the two tools is compared with the measuring stick of experimental spectroscopic information, but at least in one example where this was done, we found no clear benefit in the extra computational cost of using more sophisticated models, since CASPT2 does not prove to be superior to CASSCF and CASSCF is not superior to REC for the task of estimating energy-level distribution, including the energy of the first excited state.33 Wider and very critical reviews have also found a similar trend, when comparing effective theories vs ligand-field theory vs ab initio calculations, in the sense that neither of the approaches is a good fit for experimental results. effective theories can miss important parts of the Physics and high-level ab initio calculations tempt us to lose a critical perspective, but CASxxx methods in particular have to be considered qualitative with respect to magnetochemical properties.34
The case of HoW
Electric control of a qubit is considered to be advantageous over magnetic field (B-Field) because electric field (E-field) can be confined over a shorter length-scale. This requires the presence of strong spin-electric couplings (SECs) in the qubit medium. To seek insight into SECs phenomenon, we established a relation between spin Hamiltonian and molecular distortion as a function of the dipole moment.
Herein, we adopted a computational scheme which combines both multiconfigurational ab-initio method and an effective electrostatic approach with the objective of determining the relation between spin state and lattice coordinate. The coupling between spin states and vibrational excitations is generally accounted to be detrimental since it can contribute to decoherence, but it can also allow exploiting spin-electric couplings for coherent control. We applied this scheme to spin-qubit candidate Ho(WO) and compared the spin-electric couplings (SECs) and vibronic couplings with those already determined by perturbative approach solely using ab initio level31, 35. This scheme has already proven effective for determining key vibrations responsible for spin relaxation in molecular nano-magnets36.
This computational methodology consists of three steps to essentially access the SECs and vibronic couplings. First step is to determine the spin energy levels at multireference level (e.g.,CASSCF-SO) in crystal geometry, the obtained energy levels are employed to parameterize the ligand field Hamiltonian within the radial effective charge (REC) model37, implemented in the SIMPRE code 38. In the second step, geometry is fully optimized without restraint at DFT level to determine the vibrational frequencies and their corresponding displacement vectors. The final step consists of generating distorted geometries along the normal vectors and employing the REC model to determine the relation between spin state and lattice coordinate. This process is computationally very demanding when performed at ab-initio level, herein computational inexpensive and permit us to determine both SECs and vibronic couplings at affordable computational cast.
The spin-electric couplings are defined as shifts in transition frequency () between two spin energy levels. The relationship between applied -field and spin state is determined by noting that -field will modify the dipole-moment () of the molecule which further modifies the electric-potential. This electric-potential is stabilized by Harmonic potential of molecule, i.e. . Thus, by calculating the electric dipole moment as a function of the mode displacements, we can quantitatively extract the displacements as a function of the applied -field. Each normal mode is associated with force constant and reduced mass (yielding eigenfrequency . The electric dipole depends on the displacement of mode , and this determines the coupling of the mode to an applied -field or to incident light, that is, its infrared intensity. By linear linear combination of all normal modes, we can find an effective displacement as a function -field, e.g., . At this stage, we have determined how the spin state has evolved as a function of using the REC model described above to determine the . The results are displaced in Fig. 1 and compared with SECs determined at CASSCF-SO. We repeated this process for both DFT optimized geometry and crystallographic geometry for different applied -field (the distance between two plates where the sample is placed one could convert -field to Voltage units (V)).
The linear increment in transition frequency was observed for both optimized and crystal geometry which is in accordance with experiment and previously determined CASSCF-SO level. From Fig. 1, one could also determine that the REC model slightly overestimated the shift in transition frequency. But the overall tendency is well reproduced and satisfactory. With this linear progression spin-electric couplings constant in units of Hz/Vm are determined and placed in table 1, both methodologies resulted in same order of magnitude for SECs constant.
| Exp. | CASSCF-SO (opt.) | CASSCF-SO (crys.) | REC (opt.) | REC (crys.) | |
| SEC (Hz/Vm) | 11.4 | 2.0 | 4.5 | 6.2 | 54.5 |
The vibronic couplings detrimental for spin-dynamics are obtained by distorting the equilibrium geometry along each normal mode coordinate () and determining the evolution of crystal field parameters (CFPs). The evolution of each CFP was fitted into a second-order polynomial and evaluated versus , , allowing us to determine vibronic couplings for each normal mode. Results are displayed in Fig. 2. In particular, vibronic couplings for each normal mode are determined by first derivatives of each CFP along the respective normal mode coordinate () and finally averaging over CFPs as 39:
The obtained vibronic couplings adopting both methods, CASSCF-SO level and SIMPRE, slightly differ by 0.1, which is mainly due to not considering the second coordination sphere of HoW in the REC model. The vibrational frequency at 99.6 cm involves (HoO)w/cage tilt, vibronic coupling for this 0.087 and 0.12 cm for REC and CASSCF-SO level. Nevertheless, the overall comparison is satisfactory. This process is computationally very demanding when performed at ab-initio level, herein computational inexpensive and permit us to determine vibronic couplings at affordable computational cast.
For optimization and vibrational frequency calculations, the computing time was 160 hrs using Density function theory (DFT) implemented in Gaussian16. The spin-energy levels, using ab-initio approach (CASSCF/RASSI) method implemented in Molcas, the processing time of 10.33 hrs was spent using 4 processors in MPI parallel processing. The determination of spin energy levels for spin-electric coupling at fully ab-initio approach, a total of 13 geometries including equilibrium was calculated, a approx. of 134.33 hours processing time was used. For spin-phonon couplings, 135610.33 ( no. of mode no. of geom. processing time) 8370 hrs of computation. A similar task was performed using the REC model implemented in SIMPRE, a total processing time to determine spin-electric couplings, 13x1=13 seconds and for spin-phonon couplings, 135x6x1=810 seconds (13.5 mints) was used. The associated carbon footprint of the ab initio approach can be estimated to be in the order of 740 kg CO equivalents.40 This is comparable to km in a passenger car, or 1.3 flights (per passenger) between New York and San Francisco. Not an absurd cost, but not environmentally negligible, either. The REC approach, in contrast, would have a negligible carbon footprint in the order of 20 g CO equivalents.
Even more promising, an analytical scheme was recently demonstrated to estimate nonadiabatic coupling and state-specific energy gradient for the crystal field Hamiltonian describing lanthanide single-ion magnets. 41 Within this scheme, a single-point calculation of the desired accuracy can be employed to fine-tune parameters of a very inexpensive model, which then allows taking analytical derivatives for any desired perturbation in the molecular geometry, therefore saving hundreds of calculations that would be required to estimate the same derivatives numerically. As a result, even within frugal modelling constrictions it is possible to model spin relaxation using sophisticated nonadiabatic molecular dynamics. This clearly points towards the right direction we need to follow to keep producing good science that is compatible with the planetary boundaries: we need to know our systems well, find or develop a minimal model that recovers a good part of the relevant physics, and then, if possible employ an analytical approach to solve the problem, rather than brute-force throwing computational power into it.
Building inexpensive implementations
The ease of use of programming languages has been being increasing over time, aspect that could be identified if one searches the rise of popularity of high-level computing environments. The advantages of such type of systems is that they provide interactive exploratory environments that make language features and libraries immediately available to scientists who can use them to explore a problem domain. This contrasts with the classical edit/compile/run cycle of C or Fortran programming, which typically requires separate computation and post-processing/visualization steps. 42 Additionally, the more people use them, the more solutions to syntax errors could be found on the Internet.
If one considers the last idea, What is the effect of how we are solving our model? In this case, we explore an already published program thought to serve as a modelling tool for complex systems in the field of molecular nano-magnets and their application as probabilistic bits. 19 The fundamental reason to present this case is to emphasize that there are always different practical ways to apply what in principle is the same solution.
| Operation | File | Time in sec | |
| Matlab | Python | ||
Read input data from EXCEL file |
user configurations | 0.2346 | 0.2080 |
Read system characteristics from EXCEL file |
read_data | 0.0725 | 0.0094 |
Calculate magnetic relaxation |
mag_relaxation | 0.0021 | 0.0010 |
Calculate probabilities of each spin to flip in the 1st p-bit |
Bolztmann_distribution | 0.0010 | 0.0010 |
Iteration process (”for” loop) for the 1st p-bit |
changeable_field | 305 | 16800 |
Calculate probabilities of each spin to flip in the 2nd p-bit |
Bolztmann_distribution | 0.0473 | 0.2630 |
Iteration process (”for” loop) for the 2nd p-bit |
changeable_field | 275 | 16900 |
Average p-bits states over time |
mean_matrix_state | 0.0460 | 0.0036 |
Association analysis between both p-bits |
association | 0.1484 | 2.3946 |
Plotting results |
plotting | 0.7876 | 0.4762 |
Total time |
581 | 33700 | |
The program consists in a custom implementation of a Markov Chain Monte Carlo algorithm for each of the independent particles (in this case, effective spins ). The relative Markov chain probabilities for the spin flips between ground and excited spin states correspond to the relative Boltzmann populations of the two effective spin states , . Since each computational step has an associated natural time duration, the model allows one to follow independent time trajectories. There are three main scenarios studied using experimental data for comparison and all the details could be found in the Supporting Information File of Gutiérrez-Finol et al 19.
What is clear is that a different program, even one which follows the same algorithm, could result in a time gaining depending on the task, the size of the arrays, and complexity of the operations. Many comparison studies has been carried out regarding on the time processing between Matlab and Python; and as those investigations arise, the huge difference comes when using ”for loops” where Matlab is faster than Python. One of the main differences between M language and Python is that the first one has a strongly typed syntax. For what is known strongly typed languages can also have more efficient memory usage and processing time, because the type of each variable was already identified at compile-time; as a consequence the compiler can optimize the code saving time and being able to use the minimum amount of memory (if one declares properly the type of each variable, without the need of assuming it) and perform operations more efficiently. That being said, we’re not claiming here that the behaviour we present here is univocally associated with implementing the model in python vs in M language, since many different approaches are always possible even within a given language and algorithm.
For this research the study was carried out using a normal desk computer (Processor 11th Gen Intel(R) Core(TM) i9-11900K @ 3.50GHz, and installed memory of 16.GB with 15.8GB available) with Windows 10 Enterprise (22H2 version). What corresponds to Python (3.10 64-bit) the following modules were used: numpy, matplotlib, pandas, scipy, collections, random, math, and time. On the other hand, it was used Matlab R2023b, 64-bit.
We performed the calculations on the third case of study in a lanthanide-based, molecular spin p-bit network 19 which was the longest case in the paper. Having in mind the original case, one implemented the simulation maintaining the same conditions in both scenarios, therefore, we explored a 2-p-bit architecture where each p-bit is constituted by the collective signal of magnetic molecules. The program is divided in seven functions where each one accomplishes a specific task, two functions are reused for the calculation of the probability of each p-bit to change its state. Table 2. presents the time speed processing for each function and as it could suggest by far the largest difference is from the ”for” loops when the program iterates over each spin at each time step.
Conclusion
We presented here two particular examples illustrating an extremely common situation: one can often choose either to solve a different model or to solve it via a different implementation, and obtain significant savings in carbon footprint without making any significant sacrifices in scientific yield. Currently this is mostly overlooked, and if anything the most expensive methods tend to enjoy a higher prestige and are considered more trustworthy. Often, this means that wasteful methods allow for easier or better publishing venues. In our case, the research presented in Liu et al 31 was originally submitted with the frugal method, but during the refereeing process we switched to the wasteful method as a requisite to acceptance. As a community we need to do better, and the main factor is choosing affordable models to solve problems, with a minor but sizeable contribution of finding an inexpensive way to implement these models. Crucially, we need to beware of the Jevons’ paradox or rebound effect, meaning a more efficient method generally leads to an increased usage which often overshoots the savings. Thus, we call for frugality rather than for efficiency. Just as we consider the ethical repercussions of animal experimentation, or dealing with patient data, eventually we will need to include the risk of carbon footprint wastefulness as an ethical concern in research.
DATA AVAILABILITY
The data reported in this work and in the Supplementary Information are available from the corresponding author upon reasonable request.
Acknowledgements
A.G.A. has been supported by the COST Action MolSpin on Molecular Spintronics (Project 15128), by the European Union (EU) Programme Horizon 2020 (FATMOLS project), and by the Generalitat Valenciana (GVA) CIDEGENT/2021/018 grant. A.G.A. thanks grant PID2020-117177GB-I00 funded by MCIN/AEI/10.13039/501100011033 (both MCIN/AEI grants co-financed by FEDER funds). This study is part of the Quantum Communication programme and was supported by grant PRTR-C17.I1 funded by MCIN/AEI/10.13039/501100011033 and European Union NextGenerationEU/PRTR, and by GVA (QMol COMCUANTICA/010).
Conflict of Interest
The authors declare no competing interests.